In a survey on the gelato preferences of college students, the following data was


obtained: • 78 like mixed berry • 32 like irish cream • 57 like tiramisu • 13 like


both mixed berry and irish cream • 21 like both irish cream and tiramisu • 16


like both tiramisu and mixed berry • 5 like all three flavours above • 14 like none


of these three flavours How many students were surveyed?

1
Expert's answer
2021-12-17T04:52:58-0500
"N(B\cup C\cup T)=N(B)+N(C)+N(T)"

"-N(B\cap C)-N(B\cap T)-N(C\cap T)"

"+N(B\cap C\cap T)"

"=78+32+57-13-21-16+5"

"=122"

"N(U)=N(B\cup C\cup T)+N((B\cup C\cup T)^C)"

"=122+14=136"

136 students were surveyed.